# A Portable and Stealthy Inaudible Voice Attack Based on Acoustic Metamaterials

### v1

@@ VI-B Performance under Various Conditions
VI-B1 A1 - Range of attack on different victim devices between MetaAttack and loudspeaker array
The attack range is a critical factor affecting whether the inaudible attacks can be presented in practical scenarios; a more extended attack range can ensure the attacker’s concealment. Different victim devices can also affect the performance of MetaAttack, as the microphones and voice systems used by those have different recognition abilities. Therefore, to test the maximum attack range on various devices of MetaAttack, we let our system attack the devices (devices’ parameters are shown in Table III) from various ranges, and the experimental results are shown in Fig. 14. The outcomes demonstrate that MetaAttack exhibited optimal performance on the Apple iPhone 14 Pro, achieving a maximum attack range of 9.2 m (Success Rate = 90%). The finding suggests that MetaAttack excels particularly on Siri, potentially attributed to Siri’s superior denoising algorithm. Moreover, MetaAttack also delivered a strong performance on the other devices, with an average maximum attack range of 8.85 m (Success Rate greater than 50%). The analysis of Fig. 14 reveals that the coverage of Part A and Part B displayed in Fig. 2 is attained by all victim devices. Notably, the loudspeaker array achieved an average maximum attack range of only 2.72 m. The improved performance of MetaAttack can be attributed to the filtering and amplification capabilities of the MAM, which enables the loudspeaker array to operate at higher power and enhances its energy. These experimental results indicate that MetaAttack performs effectively across various victim devices, and its compact size further highlights its superior concealment and practicality.
@@ VI-B1 A1 - Range of attack on different victim devices between MetaAttack and loudspeaker array
Figure 14: Performance of MetaAttack on various devices, Part A and Part B refer to the quadrants in Fig 2
VI-B2 A2 - Attack range by different arrangements of loudspeaker arrays
Our simulations have shown that MetaAttack has the most effective beamforming when using 12 loudspeakers (see Section IV-C). However, experimental results indicate that the placement of the 12 loudspeakers affects the performance of MetaAttack. To identify the optimal arrangement for MetaAttack, six common settings (as shown in Fig. 17) were systematically tested in experiments, with the results shown in Fig. 17. We find that layout (a) facilitates the convergence and interference of the output sound waves from individual speaker units in a specific direction, generating a concentrated, narrow beam-like sound field. The directional nature of this sound field enhances the focusing of sound, mitigates scattering and reflection effects, and thus extends the propagation distance of the sound waves.
Figure 15: (a) Six different loudspeaker arrays. (b) Performance of loudspeaker arrays
Figure 16: Filtering performance of MetaAttack
Figure 17: Enhanced performance of MetaAttack
VI-B3 A3 - Filtering of audible leakage
To ensure the stealth of the attack, MetaAttack needs to achieve MetaAttack-side inaudibility for different inaudible commands. We use the as-k6 noise tester to test the filtering effects of MetaAttack at frequencies between 100 Hz - 4000 Hz. The experimental results are shown in Fig. 17. It has been observed that MetaAttack can effectively filter audible leakage both in front of, on sides of and behind it to levels below the “Leakage Threshold”. Moreover, the consistency between the experimental results and the filtering simulation results shown in Fig. 6 confirms the accuracy of our experimental results.
VI-B4 A4 - MetaAttack’s enhancement functionality
The enhancement functionality of MAM plays a pivotal role in enabling the reliance on 12 compact loudspeakers to exceed the long-range inaudible commands transmission achieved by large speaker array [14]. Through practical experimentation in real-world scenarios utilizing ablation techniques, we conducted comparative analyses on the impact of MAM and LAM (absent enhancement) on 12 distinct victim devices. The experimental outcomes, as shown in Fig. 17, demonstrate the substantial improvement facilitated by the enhancement feature of MAM, resulting in an average attack distance of 8.85 m, which represents a notable enhancement of 17.5% compared to LAM. Furthermore, an investigation was conducted to assess the efficacy of the inaudible commands transmitted by MAM. This was achieved through the use of the as-k6 noise tester. The results demonstrated that the energetic output of inaudible commands from MAM and LAM were 107 dB and 95 dB, respectively. These measurements serve as evidence of the superior performance capabilities of MAM.
VI-B5 A5 - Noise impact
Another crucial factor influencing the success rate of inaudible attacks is the ambient noise level. This factor is significant as noise can disrupt the accuracy of the words in the inaudible commands, leading to failure if crucial words are not correctly identified. To investigate this effect, we tested the attack range of MetaAttack across 12 distinct victim devices (consistent with A4) under varying noise conditions. The results of these experiments are illustrated in Fig. 19. In environments with noise levels approximately ranging from 55 to 65 dB (comparable to library or study areas), the maximum attack ranges of MetaAttack across different devices almost remain unaffected. However, as the noise level increases to approximately 70 to 75 dB (comparable to noisy office settings or urban traffic sounds), the maximum attack range of MetaAttack is reduced to approximately 7 - 8 m (a reduction of about 1 m). These findings suggest that the robust inaudible commands transmitted by MetaAttack demonstrate resilience against environmental noise, enabling successful inaudible attacks even in typical noise conditions.
@@ VI-C Performance in Real-world Scenarios
Figure 18: The impact of different noise levels
Figure 19: The impact of different wind speeds
@@ VI-C1 S1 - Long-range attacks between high-rise buildings
The propagation of ultrasound is influenced by the density of the medium, while changes in air density caused by wind speed impact the propagation of ultrasound. In contrast to the previous experiment A1, scenario S1 involved conducting covert and long-range attack experiments between two high-rise buildings, necessitating consideration of wind speed interference at high altitudes. In addition, the attacker is located in a public corridor with no stable power outlets nearby, and pedestrians frequently pass through the area, making it difficult to deploy large attack devices that require DC power. The efficacy of the MetaAttack in conducting attacks on target devices was evaluated under varying wind speed conditions in scenario S1, as depicted in Fig. 13a. As shown in Fig. 19, in typical wind speed environments (2 m/s - 3 m/s), the MetaAttack demonstrated a relatively stable performance, maintaining an average maximum attack distance of 8.6 m. When wind speeds exceeded 3.5 m/s, there was an impact on the attack distance. But attackers can achieve successful long-distance attacks by using weather forecasts or other means to avoid unfavorable wind conditions in advance.
VI-C2 S2 - Attacks against walking victims
Due to its portability, MetaAttack is well-suited for conducting covert attacks on walking victims in outdoor settings. In a real-world scenario illustrated in Fig. 13b, we recruited volunteers carrying various types of devices (specifically for barebones testing) for practical validation at different walking speeds. The volunteers with victimized devices maintained walking speeds of approximately 1 m/s, 1.5 m/s, corresponding to normal walking and brisk walking states, respectively. The experimental outcomes depicted in Fig. 20 demonstrate that when a victim moved at a speed of 1 m/s (average human walking pace) and was 5 m away from the MetaAttack, the success rate remained at 51.4%. It was observed that while MetaAttack maintains a consistent attack distance, the success rate of the attack significantly drops as the victim moves at a faster pace. Moreover, the efficacy of the MetaAttack varies across different victim devices. Related research [55, 56, 57] indicates that a distance of 3-4 m between individuals signifies a public distance suitable for interactions with unfamiliar individuals. Consequently, we believe that under regular walking circumstances, a trailing attack using MetaAttack beyond the three-meter threshold is unlikely to alert the victim, facilitating ease of success. To confirm the feasibility of MetaAttack, we released attack video against walking victims [32].
Figure 20: The impact of moving target’s speeds (target’s speeds are 1 m/s and 1.5 m/s)
VI-C3 S3 - Eavesdropping attacks in indoor environments
Eavesdropping is a prevalent method of information theft in commercial activities, involving covertly monitoring private conversations or communications to obtain sensitive or confidential information for unauthorized purposes. In S3 (as shown in Figure 13c), attackers can accomplish eavesdropping by sending inaudible commands to the victim’s phone to initiate a call to a number pre-set by the attacker. By leveraging the victim devices for eavesdropping, there is no necessity to physically retrieve the eavesdropping device from the premises, thereby mitigating the risk of exposure. Nonetheless, the primary challenge in this scenario lies in necessitating enhancements in character recognition accuracy. Therefore, we conducted tests to measure the word accuracy of MetaAttack in transmitting the command “call 1234567890” to various victims’ devices across different distances. As depicted in Fig. 24, the experiments indicated that MetaAttack achieved 100% word accuracy when targeting an Apple iPhone 14 Pro at a set attack distance of 8.7 m. For other devices, MetaAttack achieved 100% word accuracy within an average attack range of 8 m, demonstrating its capability to conduct eavesdropping attacks in most indoor settings.
@@ VI-C4 S4 - Effect of passing individuals on attack range
Passing individuals can significantly impact the propagation of inaudible sound waves. To investigate the influence of obstacles on the range of MetaAttack attacks at various distances, we conducted experiments using a public study room setting as an illustrative scenario. The detailed experimental parameter settings are shown in Fig. 13d. The experiment involved two passing individuals. As shown in Fig. 23, taking the iPhone 14 Pro as an example, when someone crosses MetaAttack’s attack path back and forth at a speed of 1 m/s, the system’s performance is minimally affected, allowing the attack to reach average distances of up to 7.6 m. Experimental findings indicate that passing individuals absorb and scatter ultrasonic waves, gradually attenuating the transmitted waves. But it will not significantly affect the attack range of MetaAttack.
Figure 21: Effect on attack range by passing people
Figure 22: Attack Range vs. Number of Speakers
Figure 23: The feedback accuracy of MetaAttack for different scenarios and devices
VI-C5 S5 - Feedback on the results of an inaudible attack
The feedback mechanism is crucial for inaudible attacks, as it enables the attacker to swiftly clean the attack site to evade detection by the victim. In this experiment, we evaluated the efficacy of MetaAttack’s feedback across scenarios S1 to S4. The experimental results, depicted in Fig. 23, demonstrate varying Precision and Recall values across different devices. Notably, MetaAttack exhibited enhanced performance when victim devices were Apple. Conversely, MetaAttack exhibits poorer performance in unstable scenarios (Figures 13a and 13b), indicating that external factors have a greater impact on the feedback mechanism. In such instances, the attacker must exercise timely judgment based on personal experience to avoid detection by the victim, particularly in the event of a successful and continued attack.
Figure 24: The word accuracy of MetaAttack for different devices
@@ VI-D Performance Compared to Prior Research
We made a comparison between the MetaAttack and existing primary attack devices. Theoretically, increasing the number of loudspeakers can potentially increase the attack range; however, too many loudspeakers can compromise concealment. In Fig. 23 we show the correlation between the attack range and the number of loudspeakers. An important improvement of the MetaAttack over previous studies is its invisibility and portability. As shown in Fig. 23, the MetaAttack uses only 12 speakers for its attacks and relies on only two 70 mm $\times$ 130 mm lithium batteries for power. LipRead [14] used 61 speakers for its operations and relied on power from a DC power supply. Another system called EchoAttack [9] has a limited attack range of 2.23m with 40 speakers. The enhanced DolphinAttack[15] has an extended attack range of 19.8 m, targeting the iPhone X. These enhancements are achieved through the use of 40 speakers and external accessories. However, this setup may affect the feasibility of performing covert attacks. These results show that MetaAttack is superior in terms of portability and execution of covert attacks, in contrast to previous research.
@@ VII DISCUSSION
Hardware devices for inaudible attacks. Investigating the influence of distinct attack devices on the functionality of MetaAttack system represents an interesting direction for future research. Through experimentation with a range of speaker and amplifier models, it is expected that MetaAttack system may potentially enhance its attack range efficiency while optimizing resource consumption. This aspect is worthy of further exploration and analysis in future research endeavors.
Attack through walls, windows, or curtains. The attenuation of ultrasonic signals when passing through solid objects poses a significant challenge for transmitting inaudible commands at 40 kHz. This makes it difficult for MetaAttack to effectively attack victims in enclosed environments. Future work will focus on addressing this issue.
Possible conflicts in the feedback mechanism. Considering that a few attack commands (such as ”turn on airplane mode”) may conflict with feedback mechanism, we can first confirm the position of the victim device through feedback mechanism before executing the attack.
@@ VIII COUNTERMEASURES
Previous defense strategies. The previous defense strategies against inaudible attacks can be mainly divided into two categories: forensics-based defense strategies and cancellation-based defense strategies. Forensics-based defense strategies utilize the nonlinear characteristics of inaudible commands to identify them by detecting specific differences between these commands and legitimate commands [8, 14]. In contrast, cancellation-based defense strategies by using additional ultrasonic speaker arrays to emit sound waves that cancel out the inaudible commands [58].
Challenges and Future Directions in Defending against MetaAttack Attacks. Forensics-based defense strategies can only issue warnings upon detecting an attack, while attacks from MetaAttack can often succeed within a few seconds, making it difficult for victims to timely stop ongoing attacks [58]. Although cancellation-based defense strategies can effectively prevent ongoing attacks, their effective range is limited to just a few meters. Thus, when defending MetaAttack that can conduct outdoor mobile attacks, victims must carry cumbersome equipment, such as speaker arrays and laptops, which increases inconvenience. We believe that future research should focus on developing more portable protective devices that can be carried easily or directly installed on smart devices to address the threats posed by MetaAttack.

### v2

@@ VI-B Performance under Various Conditions
VI-B1 A1 - Attack range performance between MetaAttack and loudspeaker array
Attack range is a key factor in evaluating the practicality of ultrasonic attacks—greater range offers better concealment. Device microphones and voice systems also influence MetaAttack’s performance.
We tested the maximum attack range of MetaAttack across different devices (results shown in Fig. 15). The iPhone 14 Pro performed best, achieving a 90% success rate at a distance of 9.2 meters. Other devices had an average range of 8.85 meters with success rates above 76%. MetaAttack successfully covered Regions A and B in Fig. 2 on all devices, while a regular speaker array had an average range of only 2.72 meters. The performance improvement is due to the MAM’s filtering and amplification capabilities, which significantly enhance signal strength.
@@ VI-B1 A1 - Attack range performance between MetaAttack and loudspeaker array
Figure 15: Success rate(SR) performance of MetaAttack on various devices, Part A and Part B refer to the quadrants in Fig 2.
VI-B2 A2 - Differences in speaker’s layout
Our simulation results show that the beamforming effect of MetaAttack is most significant when using 12 loudspeakers (see Section IV-C). However, we found that the speaker layout significantly impacts MetaAttack’s performance. To determine the optimal speaker layout, we simulated the sound pressure fields of six common configurations in COMSOL and systematically evaluated their RSA and SPL performance, as shown in Fig. 16. As illustrated in Fig. 16a, only layout (a) successfully forms an effective plane wave, while the others fail to do so. We attribute this to the fact that in layout (a), all speakers are positioned within the central circular opening of the metamaterial, avoiding interference from the surrounding spiral structures. Additionally, the speakers are evenly spaced and symmetrically arranged, which promotes coherent wavefront formation and facilitates the generation of a desirable plane wave. In contrast, the other layouts either suffer from obstruction by the spiral structures or lack consistent spacing and symmetry, which hinders effective wavefront shaping. Fig. 16b further confirms that layout (a) achieves the best performance in both SPL and RSA, leading us to select it as the final design for the system.
Figure 16: (a) Six speaker arrays with their sound fields. (b) Performance of speaker arrays (tests include arrays’ RSA and SPL at Transmitter(TX).
Figure 17: (a) Leakage SPL comparison at TX (top) and its relation to the leakage threshold (bottom). (b) Volunteers’ perception of the leakage.
Figure 18: (a) SPL comparison of inaudible attacks with and without MAM. (b) RSA performance comparison.
VI-B3 A3 - Noise impact
Another factor affecting the success rate of inaudible attacks is environmental noise, which can interfere with the accuracy of command recognition. We tested MetaAttack’s RSA on 12 devices under varying noise conditions. The results, shown in Fig. 20, indicate that only when the noise level increases to 70-75 dB (similar to noisy office or urban traffic sounds) does the attack range decrease to 7-8 meters (a reduction of about 1 meter). This suggests that the strong beamforming effect of MetaAttack allows it to successfully perform inaudible attacks even in typical noisy environments.
VI-B4 A4 - Leakage filtering performance
To ensure the stealth of the attack, MetaAttack needs to effectively filter out audible leakage. In a laboratory environment with a background noise level of 30 dB, we played single-tone signals ranging from 100 to 4000 Hz at 60 dB SPL and used the as-k6 noise tester to evaluate the filtering performance of MetaAttack. As shown in Fig. 17a, MetaAttack significantly attenuates sound waves from the front, sides, and rear, reducing audible leakage to below the Leakage Threshold.
To verify whether the sound filtered by MetaAttack is imperceptible to the human ear, we invited 42 volunteers to participate in the experiment on a voluntary basis. The volunteers were evenly divided into three age groups (14 volunteers per group): 8–12 years old (children, more sensitive to sound), 13–18 years old (teenagers, broader hearing range), and 19–30 years old (young adults, peak auditory resolution), with a balanced gender distribution within each group. The experiment was conducted in an environment with a background noise level of 30 dB. Volunteers stood at various distances and attempted to identify whether there was any audible leakage at four levels: inaudible, mild noise, indistinct and intelligible [19, 12, 20, 64].
As shown in Fig. 17b, the experimental results indicate that only within a short distance of 0.1 to 0.5 meters could a few volunteers with more sensitive hearing perceive faint or indistinct sounds. We speculate that this may be due to slight audible leakage caused by nonlinear acoustic effects during air transmission. However, our attack range far exceeds this distance (reaching up to 8.85 meters), and the background noise level in the experimental environment was only 30 dB, which is extremely quiet. In real-world environments, such weak audio leakage is typically masked by higher background noise. Therefore, we believe that MetaAttack demonstrates strong imperceptibility in practical attack scenarios.
VI-B5 A5 - Contribution of metamaterials
The enhancement functionality of MAM plays a key role in enabling 12 compact loudspeakers to surpass large speaker arrays in achieving long-range inaudible command transmission [19]. Through experiments conducted in real-world scenarios using ablation techniques, we compared the RSA of MAM, LAM and speaker array without metamaterial on 12 distinct target devices, and analyzed the SPL at various distances.
As shown in Fig. 18a, the results show that MAM (8.85 m) improves the average attack distance by 14.2% compared to LAM (7.75 m) and by 225% compared to the speaker array (2.72 m). Compared to LAM (95 dB), MAM (107 dB) increases the output SPL by 12.6%, and by 28.9% compared to the speaker array (83 dB). These results highlight the two key functions of MAM—audible leakage suppression and energy amplification—as essential to enhancing speaker array performance.
@@ VI-C Performance in Real-world Scenarios
Figure 19: Impact of noise levels.
Figure 20: Impact of wind speeds.
@@ VI-C1 S1 - Long-range attacks between high-rise buildings
As shown in Fig. 14a, the experiment in Scenario S1 was conducted between two high-rise buildings, taking into account high-altitude wind interference. The attacker was located in a public corridor with no power supply and frequent pedestrian traffic, making it impractical to use large equipment. The experimental results (Fig. 20) show that MetaAttack maintained a stable attack range of 8.6 meters at wind speeds of 2–3 m/s, while wind speeds above 3.5 m/s slightly affected the attack distance. However, attackers can rely on weather forecasts to avoid unfavorable wind conditions and ensure reliable long-range attacks.
VI-C2 S2 - Attack against outdoor targets
Due to its portability, MetaAttack is well-suited for conducting covert attacks on moving targets in outdoor environments. To validate this, we recruited volunteers carrying different types of devices in the real-world scenario shown in Fig. 14b. The volunteers walked at speeds of 1 - 1.5 m/s, corresponding to normal walking and brisk walking, respectively, with the devices either held in hand or placed in pockets.
The experimental results in Fig. 22 show that when the target moves at a speed of 1 m/s and holds the device, MetaAttack achieves its optimal average RSA of 6.2 meters. As the target’s speed increases or the device is placed in a pocket, the attack success rate decreases. This is due to slight shaking of the device during movement, making it difficult for the system to align, and physical obstructions like a pocket weakening the signal. However, even in the harshest conditions (1.5 m/s speed with the device in the pocket), MetaAttack still achieves an average RSA of 4.1 meters. As mentioned in Section II-B, a distance beyond 3 meters is considered a safe range unlikely to alert the target. Therefore, we believe that MetaAttack can perform attacks without alerting the target. To validate the feasibility of MetaAttack, we released an attack video targeting walking individuals [37].
Figure 21: Impact of outdoor target.
Figure 22: System’s Word Acc. (WA)
VI-C3 S3 - Eavesdropping attacks on sensitive information
Eavesdropping is a common method of information theft in commercial activities, where attackers send inaudible commands to manipulate the target phone to dial a preset number, as shown in Fig. 14c. The main challenge in this scenario is improving character recognition accuracy. We evaluated the performance of MetaAttack in transmitting the command call 1234567890 at various distances. Results show that MetaAttack achieved 100% recognition accuracy at a distance of 8.85 meters on an Apple iPhone 14 Pro, and similar accuracy on other devices within an average range of 8 meters. This demonstrates that the metamaterial-based beamforming design of MetaAttack enables efficient and precise signal transmission.
@@ VI-C4 S4 - Effect of passing individuals on attack range
Passing pedestrians may affect the propagation of inaudible sound waves. To assess their impact on the attack range of MetaAttack, we conducted experiments in a public study room (parameters shown in Fig. 14d). As shown in Fig. 24, when a person walks back and forth along the attack path at a speed of 1 m/s (using the iPhone 14 Pro as an example), the attack distance still reaches 7.6 meters with almost no performance degradation. Although pedestrians may absorb and scatter part of the ultrasonic signal, the beamforming design of MetaAttack ensures that the sound waves remain highly directional and focused, allowing the signal to be transmitted effectively without being significantly dispersed, thus maintaining the effectiveness of the attack.
Figure 23: Effect by passing people
Figure 24: Performance compared to prior research (attack range vs. number of speakers)
Figure 25: The feedback accuracy of MetaAttack for different scenarios
VI-C5 S5 - Feedback mechanism performance
The feedback mechanism is crucial for inaudible attacks, as it enables the attacker to swiftly clean the attack site to evade detection by the target. In this experiment, we evaluated the efficacy of MetaAttack’s feedback across scenarios S1 to S4. As shown in Fig. 25, although the Precision and Recall vary slightly across different devices, they consistently remain above 97%, demonstrating high feedback accuracy. Notably, MetaAttack performs relatively worse in unstable scenarios (Fig. 14a and Fig. 14b), indicating that the feedback mechanism is more susceptible to external environmental interference. In such instances, the attacker must exercise timely judgment based on personal experience to avoid detection by the target, particularly in the event of a successful and continued attack.
@@ VI-D Performance Compared to Prior Research
We made a comparison between the MetaAttack and existing primary attack devices. Theoretically, increasing the number of loudspeakers can potentially increase the attack range; however, too many loudspeakers can compromise concealment. In Fig. 24 we show the correlation between the attack range and the number of loudspeakers. An important improvement of the MetaAttack over previous studies is its invisibility and portability. As shown in Fig. 24, the MetaAttack uses only 12 speakers for its attacks and relies on only two 70 mm $\times$ 130 mm lithium batteries for power. LipRead [19] used 61 speakers for its operations and relied on power from a DC power supply. Another system called EchoAttack [13] has a limited attack range of 2.23m with 40 speakers.
The enhanced DolphinAttack[20] has an extended attack range of 19.8 m, targeting the iPhone X. These enhancements are achieved through the use of 40 speakers and external accessories. However, this setup may affect the feasibility of performing covert attacks. These results show that MetaAttack is superior in terms of portability and execution of covert attacks, in contrast to previous research.
@@ VII DISCUSSION
Hardware devices for inaudible attacks. Exploring the impact of different attack devices on the functionality of the MetaAttack system is an interesting direction for future research. By experimenting with various speaker and amplifier models, the MetaAttack system is expected to improve attack range efficiency while optimizing resource consumption.
Attack through walls, windows, or curtains. The attenuation of ultrasonic signals when passing through solid objects poses a significant challenge for transmitting inaudible commands at 40 kHz. This makes it difficult for MetaAttack to effectively attack targets in enclosed environments. Future work will focus on addressing this issue.
Audible leakage in the air propagation process. The audible leakage during airborne transmission caused by nonlinear acoustics remains an issue to be addressed. Although MetaAttack has compressed the audible leakage into a narrow range of about 12°, where it can only be detected if the target’s ear is within this range, the problem is not yet fully resolved. We plan to thoroughly address this limitation in future work.
Optimization of Metamaterials. Currently, MetaAttack has achieved the longest possible attack range under its existing structure. We believe that further extending the attack distance hinges on enhancing the filtering performance of the metamaterials and improving their beamforming capability at lower carrier frequencies (e.g., around 25 kHz). Compared to high-frequency ultrasound, low-frequency ultrasound maintains inaudibility while offering better propagation characteristics, making it more promising for long-range transmission. In addition, due to the constraints of the fixed LAM filtering structure, the current optimization of speaker parameters (such as quantity, arrangement, and carrier frequency) may only yield locally optimal results. In future work, we plan to explore more advanced metamaterial designs and jointly optimize them with speaker parameters to enable more stable and longer-range attacks.
Possible conflicts in the feedback mechanism. Considering that a few attack commands (such as ”turn on airplane mode”) may conflict with feedback mechanism, we can first confirm the position of the target device through feedback mechanism before executing the attack.
@@ VIII COUNTERMEASURES
Previous defense strategies. Defense strategies against inaudible attacks fall into two main categories: forensics-based and cancellation-based. Forensics-based defenses detect attacks by analyzing signal characteristics through methods like frequency analysis or deep learning. For example, EarArray [62] analyzes timing differences across microphones; Li et al. [65] extract spatial features and apply deep learning to distinguish human speech from machine-generated signals; NormDetect [63] identifies missing features in attack audio; and LipRead [19] analyzes power, autocorrelation, and low-frequency amplitude deviations. Cancellation-based defenses use ultrasonic speaker arrays to emit sound waves that neutralize attack signals [66].
Challenges and future directions in defending against MetaAttack attacks. Forensics-based defenses are sensitive to microphone variations, leading to unstable results [66]. We replicated LipRead [19] in a quiet (40 dB) lab and tested 30 rounds on various devices. iPhone 13 Pro Max, iPhone 13, iPhone 12, Galaxy S7, and nova 7 SE 5G showed over 90% success, while iPhone 14 Pro, Redmi K50 Ultra, Pixel 3, and Pixel 2 were below 63.3%, confirming poor cross-device consistency. These defenses also disable voice assistants only after detecting attacks, disrupting normal use. Cancellation-based methods are more effective but rely on bulky equipment and have limited range, making them impractical for mobile attacks like MetaAttack. Future defenses should focus on lightweight, externally integrated mechanisms that block attack signals before they reach the microphone.